urgent..
working together, alice and betty can do a certain job in 4 1/3 days. but alice felt ill after 2 days of working and betty finished the job continuing to work alone in 6 3/4 more days. how long would it take each to do the job if each of them worked alone?
thnx for the help

Answer by ptaylor(2198)

Alice and Betty works at the rate of 1/(4 1/3) or 1/(13/3)=3/13 job per day
Working together, in 2 days they do(3/13)*2=6/13 of the job, Leaving 7/13 of the job yet to be done. Now we are told that Sally finishes the job in 6 3/4 days which means that she works at the rate of (7/13)/(6 3/4) jobs per day.
(7/13)/(6 3/4)=(7/13)/(27/4)=(7/13)*(4/27)=28/351 jobs per day
Now, we have determined that Sally works at the rate of 28/351 jobs per day
If we let x=Alices rate of work , we have the following equation to solve:
x + 28/351=3/13=81/351 subtract 28/351 from each side
x=81/351 - 28/351=53/351 jobs per day---Alice's rate of work
Now if Sally works at the rate of 28/351 jobs per day, in zs(z sally) days, she does 1 job. so:
(28/351)*zs=1
zs=351/28 =~12.5 days--------time it takes Sally working alone
Now for Alice:
(53/351)*za=1
za=351/53 =~6.6 days-----------time it takes Alice working alone
CK
28/351 + 53/351= 3/13
81/351=3/13
3/13=3/13
Hope that you can follow what I've done---ptaylor